# Problem Solving Through Human-AI Preference-Based Cooperation

### v3

@@ Abstract
While there is a widespread belief that artificial general intelligence (AGI) – or even superhuman AI – is imminent, complex problems in expert domains are far from being solved. We argue that such problems require human-AI cooperation and that the current state of the art in generative AI is unable to play the role of a reliable partner due to a multitude of shortcomings, including inability to keep track of a complex solution artifact (e.g., a software program), limited support for versatile human preference expression and lack of adapting to human preference in an interactive setting. To address these challenges, we propose HAI-Co2, a novel human-AI co-construction framework. We formalize HAI-Co2 and discuss the difficult open research problems that it faces. Finally, we present a case study of HAI-Co2 and demonstrate its efficacy compared to monolithic generative AI models.
@@ 1 Introduction
Despite the impressive achievements of generative AI spearheaded by Large Language Models (LLMs), Vision Language Models and code models (Lozhkov et al., 2024; Wang et al., 2021), multiple recent investigations have pointed out their lack of competence in dealing with complex generation problems that require intricate planning (Kambhampati et al., 2024) or task adherence while keeping track of multiple constraints (Xie et al., 2024). A broad class of such complex problems requires active human participation. Therefore, although the recent focus in generative AI has mostly been on complete automation, we believe that human-AI cooperation is a more promising approach for complex problems of this kind.
To effectively support human-AI cooperation, we draw inspiration from how humans collectively address complex problems to devise solutions: Humans often solve complex problems by iteratively co-constructing a solution step by step, revising and refining draft solutions while transitioning between different levels of abstraction, and exchanging information about preferences and potential improvements in natural language (Damşa, 2013). Our position is that this type of human-human cooperation is a promising template for the cooperation of humans and AI agents. On this basis, we propose Human-AI Co-Construction (HAI-Co2), a novel framework for human-AI cooperative problem solving that builds on preference-based learning and search methodology and relies on natural language to facilitate interaction. The problem solving process is conceived as a process of systematic search in a construction space $\mathcal{X}$ of candidate solutions, i.e., as a co-constructive process, in which candidate solutions are modified step by step until a sufficiently good solution has been found.
Figure 1: Existing generative AI lacks proficiency in key aspects of co-construction of solutions to complex problems. We give a code synthesis example. (a) GPT-4 Turbo fails to follow preferences explicitly stated by the human expert. (b) Due to the lack of a persistent object representation, a modification request targeted toward one feature of the desired solution leads to the unwanted (and erroneous) modification of another feature. (c) Human expert modifies the generated code directly to remove inline assumptions and introduce general variables; such active participation is not demarcated and recorded by the AI and there is no facility to extract the implicit preference and follow it elsewhere.
What are co-construction problems? The broad class of problems that we seek to address in this paper are primarily centered around domain-specific expert applications where the task is to construct a solution that meets specific requirements, e.g., a computer program in software engineering or a machine learning pipeline in automated machine learning (AutoML). We choose the term co-construction because of the necessity for the AI agent to work closely with human experts.111 Not all problems require co-construction; some have unique solutions independent of human feedback. As AI models become more capable, they will be able to solve such problems, which currently require co-construction due to limited AI capability, without help from the human expert since they can better utilize prior knowledge and task specifications. However, in many cases, task descriptions remain inherently incomplete due to either ambiguous specifications or evolving expert preferences, necessitating co-construction with human input regardless of model capability. First, capabilities of current AI systems are limited and cannot fully replace human expertise in these tasks, e.g., due to insufficient knowledge and reasoning capabilities, lack of trustworthiness, and bias. Scaling generative AI systems, particularly language models, has demonstrated improved performance across a range of tasks, e.g., mathword problems and commonsense reasoning. However, even the most powerful LLMs show a lack of robustness under different semantic perturbations that would not have fooled an otherwise robust reasoning machine (Li et al., 2024). While it is unwise to rule out future improvements, their current limitation certainly calls for interventions beyond scaling. Second, some expert tasks (e.g., designing websites and building data analysis and visualization software) require irreplaceable human supervision; solutions to be constructed in these problems are defined based on personalized human preferences. While the co-constructed solution must fulfill a set of objective correctness criteria that can be verified using symbolic tools without any human intervention, the human expert imposes a broad set of subjective criteria as well. For example, in a software development setup, the final piece of software must be executable, secure, and bug-free; at the same time, every expert developer has their own style – of coding, commenting, modularization, etc. – that should be followed as well. Similarly, an ML pipeline developer may want a suboptimal classifier to save computational cost and go for optimality in a later stage of development. This implies that the AI agent must be able to follow and adapt to the personalized preferences of the human expert (which often are not stable, but evolve in the course of the problem solving process).
Since we see expert domains as the primary setting in which co-construction excels at solving problems, we use “expert” and “human” interchangeably in this paper.
Is the current state of generative AI enough? Multiple prior investigations have laid out the inherent shortcomings of present-day generative AI that limit its applicability to the type of problems we are addressing in this paper. In an example case study tailored towards code generation presented in Figure 1, we demonstrate some of the major bottlenecks of GPT-4 Turbo, a recent update [^2] We use the gpt-4-1106-preview instance: https://openai.com/index/new-models-and-developer-products-announced-at-devday/ over the original GPT-4 model (OpenAI, 2024). In Figure 1 (a), GPT-4 Turbo ignores the human expert’s explicit instructions to generate a complete Python code with the required module specifications, echoing Xie et al. (2024)’s observation that current language-based AI agents lack task adherence. After repeated prompting with partial code snippets, the process produces a complete – albeit faulty – code. This limitation becomes even more irreconcilable as more varied and realistic expressions of human preferences are taken into account – for the human expert to contribute productively, one must allow preferences expressed via explicit instructions, binary choice, ranking, etc. Current generative AI solutions do not facilitate such multi-modal preference incorporation. Figure 1 (b) shows unreliable debugging attempts. Specifically, the LLM performs unrelated (and faulty) edits to address a bug and even introduces new errors. This demonstrates that existing LLMs fail to handle complex, modular software code (Jiang et al., 2024). The common practice is that the human (as a knowledgeable expert who keeps track of overall context) identifies faulty output and repeatedly prompts the model to guide it to the correct generation – this is implicitly adopting a co-construction paradigm. However, Figure 1 (c) shows that current modes of human-AI interaction cannot unleash the full potential of co-construction – direct modification of the co-constructed candidate solution by the human expert does not bear any special significance to the LLM, and it treats it as just another context. There is no explicit mechanism for the AI to learn implicit preferences expressed by the human through active participation.
While these examples are focused on code synthesis, there is evidence of similar shortcomings in other domains (Kambhampati et al., 2024; Lecler et al., 2023; Almarie et al., 2023). Carroll et al. (2019), for example, demonstrate the necessity of incorporating explicit “human-awareness” in a version of the collaborative game Overcooked, providing evidence that agents fail to coordinate with human subjects without such awareness. Code synthesis in particular – and the experience from day-to-day use of generative AI for solving complex problems in general – points toward co-construction as a naturally evolved problem solving paradigm where the human expert tries to search for the optimal solution by interacting with the AI. However, the current state of generative AI hinders its role as a reliable partner in successful co-construction. This is because the “one-directional” interaction between human and AI typical of how AI agents are used today often fails to steer the co-construction towards a solution that satisfies the user’s constraints.
Our contribution. In this paper, we formalize co-constructive problem solving and thereby aim to address important limitations of current generative AI models. We present HAI-Co2, a conceptual framework with three fundamental properties that facilitate human-AI co-construction. First, it introduces multiple levels of abstractions to the candidate solution, providing a seamless interface for the human expert and the AI agent to modify and keep track of the complex, modular, co-constructed candidate solution. Second, HAI-Co2 allows multi-modal preference input from the human expert, with natural language as the central mediator to capture information-rich guidance signals, along with other forms of active expert participation, such as categorical choice-based preference. Finally, HAI-Co2 introduces a search-based methodology of co-construction where the candidate solution (represented on multiple levels of abstraction) is iteratively revised to maximize the perceived utility modeled from the preference input. While multiple components of HAI-Co2 have been explored in prior research independently across different domains, we are the first to bring them together under a unified conceptual umbrella and to show that jointly with current generative AI, they have the potential to address multiple major challenges in solving expert domain problems. We showcase the efficacy of HAI-Co2 using an example implementation for the expert domain of software engineering.
The rest of the paper is organized as follows. After a discussion of relevant literature in Section 2, we introduce the framework of HAI-Co2 in Section 3 and outline open research challenges implied by this framework in Section 4. Section 5 illustrates HAI-Co2 in a case study, prior to concluding the paper in Section 6.
2 Related work
We group relevant prior work into five major strands: reinforcement learning from human feedback (RLHF), assistance games, search- and evolution-driven construction, LLM agents for complex problem solving, and persistent solution spaces for iterative construction. We discuss how these approaches attempt to address expert-AI co-construction. However, as we will see in Section 4, they fail to comprehensively tackle these challenges.
Reinforcement Learning from Human Feedback. RLHF focuses on learning a policy preferred by humans, most commonly relying on comparisons between candidate solutions (Kaufmann et al., 2024). The goal is to learn a policy that maximizes a reward or utility function that is consistent with the human feedback. Originating in classical reinforcement learning domains such as games and continuous control (Christiano et al., 2017), RLHF has been extended to a variety of domains, most notably fine-tuning generative models such as LLMs (Stiennon et al., 2020; Ouyang et al., 2022), eventually leading to the development of AI models that can generate human-preferred responses in natural language such as ChatGPT.
Assistance Games. Another approach rooted in reinforcement learning is assistance games (Shah et al., 2021; Laidlaw et al., 2024), originally introduced as cooperative inverse reinforcement learning (Hadfield-Menell et al., 2016). As a framework for cooperative decision-making between an AI agent and a human, assistance games model human-AI interaction as a cooperative game with the agent and the human as players. The AI agent aims to maximize the human’s utility function without explicit knowledge of it. It learns this function by observing the human’s actions and receiving feedback on its own actions, which may include queries to the human.
The process of co-construction in the sense of HAI-Co2 could in principle be seen as an assistive sequential decision-making problem, too, with the expert’s and the agent’s actions being steps taken to construct the solution. However, in practice, casting co-constructive problem solving as an assistance game does not seem particularly useful to us. Roughly speaking, for HAI-Co2 we consider co-constructive problem solving as a search rather than a reinforcement learning task. While both tasks are related, RL is a very general setting that comes with additional challenges, notably the learning of a model through exploration (in a non-deterministic environment with unknown effects of actions). Indeed, the price that RL pays for its generality is a high conceptual and computational complexity. This problem applies in particular to assistance games, which, despite their nice theoretical properties (Shah et al., 2021), have not yet been widely applied to real-world problems. [^3] Laidlaw et al. (2024) present a first step towards scalably solving assistance games in more complex environments. This is early work, though, and we believe that the conceptual and computational benefits of the search-based formulation justify a more specialized framework for this highly relevant use case. We hope that our framework can provide a more practical approach to human-AI co-construction by focusing on the specific challenges of solution construction with a search process guided by the extensive prior knowledge of pretrained language models.
Learning from natural language interactions. Natural language interaction between the expert and the AI agent is central to HAI-Co2. A popular approach towards facilitating human-AI interactive problem solving involves training the agent to follow instructions in natural language (Branavan et al., 2009; Tellex et al., 2011). This paradigm of learning to follow instructions has found attention in the LLM era as well (Wei et al., 2022). However, directly mapping language-specified goals to actions has limited applicability to novel tasks. Alternatively, learning rewards from natural language interaction to successfully align the AI agent with the human user has been explored (Fu et al., 2019; Sumers et al., 2021) – instead of learning to map language-defined goals to actions directly, they seek to learn the reward function from the language-defined goals that can be generalized to novel tasks. The findings of Sumers et al. (2022) suggest that while instructions typically perform well in low-autonomy settings, high-autonomy regimes favor the reward-learning paradigm. Instead of learning a language-conditioned policy or reward, it is also possible to use conversational cues as rewards themselves (Jaques et al., 2020), which can be combined with the other approaches discussed here. Yet another approach, deployed in OpenAI’s ChatGPT, is to enable the language model to save information about the user and their interactions with the model as natural-language ‘memories’, which may include information about the user’s preferences. [^4] https://openai.com/index/memory-and-new-controls-for-chatgpt/ Most of these approaches do not consider the need to actively elicit and co-construct the expert’s preferences, a key aspect of HAI-Co2. This necessity is supported by Co-Reyes et al. (2019) and Lin et al. (2022), who identify that inferring the correct behavior or reward from a single utterance can become non-trivial given the multidimensionality of language. Querying the human user and estimating their preferences in an interactive setup is also a key component of Peng et al. (2024)’s framework. As a step in this direction, Li et al. (2023a) use active elicitation to strengthen preference understanding – the AI agent is trained to elicit and infer human preferences by actively interacting with the user. This is crucial prior work for an implementation of HAI-Co2, forming an important pillar of future research under the abstract umbrella it provides.
Search- and Evolution-Driven Construction. Our framework emphasizes iterative search within the construction space, a process akin to evolutionary optimization, which iteratively generates and evaluates candidate solutions (Bäck, 1996). This evolution can be viewed as a form of search-based construction. Interactive evolutionary computation, a preference-based extension, is particularly relevant to our work as it involves human evaluation of candidate solutions (Takagi, 2001; Wang & Pei, 2024). For example, these methods have been applied to search-based procedural content generation in video games (Togelius et al., 2011). Our approach differs in the core approach to the search process: Traditional evolutionary methods maintain a population of candidate solutions and generate new ones through mutation and recombination. In contrast, in our framework, each iteration ends with a single candidate solution that is then the basis for the next iteration. In addition, we leverage the extensive prior knowledge of pretrained language models to guide the search and use natural language communication to facilitate cooperation between the AI agent and the human expert.
LLM agents for complex problem solving. The rapid increase in the capabilities of LLMs has triggered multiple recent efforts to integrate them at the core of autonomous agents that interact with the environment, plan, and act to solve complex problems (Wang et al., 2024). Typical approaches adopt integrating different tool-usage capabilities into LLMs via efficient prompting, often with multimodal capabilities (Chen et al., 2023). A single agent is often insufficient to solve complex problems; thus, multiple agents with different capabilities have to be integrated. Recent efforts in LLM-based multi-agent systems seek to mimic such cooperative problem solving by role-playing LLMs via in-context examples (Li et al., 2023b) or fine-tuning (Juneja et al., 2024). Despite some promising achievements, such agentic ecosystems are inherently limited by the constituent LLMs’ inability to plan and execute tasks (Xie et al., 2024; Kambhampati et al., 2024). These frameworks of LLM-based autonomous agents are largely designed for autonomous problem solving, not for human-AI co-construction as we envision. Recently, arguments in favor of strategically allocating tasks between humans and LLM-based agents to exploit their distinct strengths have been put forward (He et al., 2024), which aligns with our approach of leveraging the strengths of both humans and AI agents in co-construction. An alternative view of our framework is hence an extension of LLM-based multi-agent systems with a human agent as a key component, focusing on the co-construction of solutions.
Figure 2: Illustration of the hierarchy of construction spaces in HAI-Co2. Each point ${\bm{x}}$ symbolizes a candidate solution (on a certain level of abstraction), e.g., a software program. The topology of the space is specified by a suitable neighborhood structure (as illustrated for point ${\bm{x}}$). Each point is associated with a latent utility $u$, possibly multi-dimensional and comprised of local utilities $u_{1},\ldots,u_{K}$, and preferential information (e.g., ${\bm{x}}\succ{\bm{x}}^{\prime}$: solution ${\bm{x}}$ is better than ${\bm{x}}^{\prime}$) that provides information about promising regions in the space. The relationship between the different abstraction levels is specified by the abstraction mappings $f_{j}$ resp. the (inverse) refinement mappings $f^{(-1)}_{j}$.

### v4

# Position paper

@@ Abstract
While there is a widespread belief that artificial general intelligence (AGI) – or even superhuman AI – is imminent, complex problems in expert domains are far from being solved. We argue that such problems require human-AI cooperation and that the current state of the art in generative AI is unable to play the role of a reliable partner due to a multitude of shortcomings, including difficulty to keep track of a complex solution artifact (e.g., a software program), limited support for versatile human preference expression and lack of adapting to human preference in an interactive setting. To address these challenges, we propose HAI-Co2, a novel human-AI co-construction framework. We take first steps towards a formalization of HAI-Co2 and discuss the difficult open research problems that it faces.
@@ 1 Introduction
Despite the impressive advances of generative AI (Cao et al., 2025), especially for natural language (large language models), vision (vision language models) and code (code models), recent investigations have pointed out a lack of competence in dealing with complex generation problems that require intricate planning (Kambhampati et al., 2024) and task adherence while keeping track of multiple constraints (Xie et al., 2024). A broad class of such complex problems, especially complex problems in expert domains, requires active human participation. Therefore, although the recent focus in generative AI has mostly been on complete automation (Hong et al., 2024; Brown et al., 2020), we believe that human-AI cooperation is a more promising approach.
To address complex problems of this kind, we propose Human-AI Co-Construction (HAI-Co2), a novel framework for human-AI cooperative problem solving that builds on preference-based learning and search methodology and relies on natural language to facilitate interaction. In this article, we first give a general introduction to HAI-Co2 (Section 2), including its ethical challenges, and take first steps towards a formalization (Section 3). We outline open research challenges posed by HAI-Co2 in Section 4. Section 5 discusses related work. Section 6 concludes the paper.
@@ 2 HAI-Co2: Human-AI co-construction through preference-based search
In this article, we propose Human-AI Co-Construction (HAI-Co2), a novel framework for human-AI cooperative problem solving. The four defining characteristics of HAI-Co2 are (i) solution of complex problems in expert domains that require active human participation, (ii) co-construction of the solution by human and AI agent, (iii) co-construction of the objective by human and AI agent by means of preference learning and (iv) the use of natural language as the main communication medium, which makes it possible for human and AI agent to be equal partners, with complementary strengths, in the co-construction. We now describe these four characteristics in more detail.
First, we target applications in expert domains where the task is to construct a solution to a complex problem. Since expert domains are our focus, we use “expert” and “human” interchangeably in this article. Examples of complex problems in expert domains include writing a computer program in software engineering; constructing a machine learning pipeline in automated machine learning (AutoML); writing a related work section in scientific research; and developing a formalization of a problem described in natural language in mathematics.
Second, we conceive of problem solving as co-construction of the solution to the complex problem by a human and an AI agent or – more generally – by a team of humans and agents. As detailed in Section 3, the problem solving process is formalized as a process of systematic search in a construction space $\mathcal{X}$ of candidate solutions on several hierarchical levels of abstraction. In this co-constructive process, candidate solutions are modified step by step until a sufficiently good solution has been found.
We draw inspiration from our understanding of how humans collectively devise solutions to complex problems. Humans often tackle such problems by iteratively co-constructing a solution step by step, revising and refining draft solutions while transitioning between different levels of abstraction and exchanging information about preferences and potential improvements in natural language. Our primary motivation for HAI-Co2 comes from natural language processing and computer science; see Section 5 for a brief discussion of related fields that have conducted extensive research on human-human cooperation on solving tasks.
Human-human cooperation would make less sense as a promising template for human-AI cooperation if current AI systems could solve complex expert-domain problems on their own. However, current AI capabilities are limited for complex expert-domain problems, e.g., due to insufficient knowledge and reasoning capabilities, bias and lack of trustworthiness. Scaling generative AI systems, particularly language models, has demonstrated improved performance across a range of tasks, e.g., math word problems and commonsense reasoning. However, even the most powerful LLMs show a lack of robustness under different semantic perturbations that would not have fooled an otherwise robust reasoner (Li et al., 2024). While it is unwise to rule out future improvements, their current limitations call for interventions beyond scaling. Thus, in order to be able to effectively solve complex expert-domain problems, we believe it is necessary for the AI agent to work closely with human experts.
Third, cooperative problem solving often involves the co-construction of the objective – or objective co-construction – alongside the co-construction of the solution itself. As part of the co-construction process, the requirements for the solution are often changed and refined as the collaborators understand the details of the complex problem better and revise their initial assumptions. As we will see in Section 3, we formalize this process of objective co-construction through interactive preference learning: we define a utility function over the construction space that reflects preferences of user and agent, i.e., which artifacts are better and which are worse candidate solutions. Thus, the objective is encoded in a preference model. Or, stated differently, the objective is first described informally on a general level – e.g., write a computer program performing a particular task – and then formalized in terms of the preference model.
One inspiration for making objective co-construction part of our framework for solution co-construction comes from the field of multi-criteria decision aiding (MCDA), a branch of operations research Roy (2000, 1996). “MCDA underlines the ’aiding’ in a process involving the DMs [decision makers] in the co-construction of their preferences …  It assumes that preferences of the DM with respect to considered alternatives do not pre-exist in the DM’s mind.” (Hüllermeier and Slowinski, 2024a) MCDA (see also Hüllermeier and Slowinski (2024b)) makes assumptions that are close to objective co-construction; in particular, the user’s objective is only partly determined in the beginning and further developed in the course of the decision aiding process.
It is important to note that the two types of co-construction in HAI-Co2 – solution co-construction and objective co-construction – are quite different. Solutions are artifacts whereas objectives are encoded as preference models. The solution is co-constructed through search in the construction space whereas the objective is co-constructed through preference learning. Still, at the highest level, both solutions and objectives are the result of a cooperation of human and agent, with each being a contributor even though their contributions may differ in nature and scope.
Fourth, in HAI-Co2, natural language is the main communication medium. This makes it possible for human and AI agent to be equal partners, with complementary strengths, in the co-construction process. In our view, language-based communication is a key enabler of an equal partnership. The ability to express oneself fully and on the same level is a prerequisite for making equal contributions to problem solving. Previous communication technology was a limiting factor in this regard: only with the advent of LLMs do we now have AI agents available that comprehend and generate natural language at a human level of capability. Such human-level capabilities are required for the complex communication needs that occur during cooperation on complex expert-domain problems.
There are certainly problem-solving scenarios in which the human manages the process and the agent’s role is reduced to handling low-level tasks (e.g., “tool” tasks like internet search or copy-editing) – or, conversely, scenarios where the human’s role is reduced to providing input when prompted by the agent (e.g., in active learning). In contrast, the type of human-AI cooperation scenario we are interested in is one in which the two partners are equal.
Equality here does not mean identical roles. On the contrary, the roles are complementary: each partner has skills or knowledge that the other lacks. An example for complementarity is that the human may understand the context better in which the complex problem arises (e.g., the requirements and personalized preferences of human stakeholders) whereas the AI agent may be able to more efficiently access vast information resources and make more effective use of tools such as compilers and unit testing. Just as human-human cooperation excels at problem solving if the collaborators complement each other, so is human-AI cooperation most beneficial if each partner can contribute their unique strengths. This aligns with recent evidence showing that human-AI cooperation alone does not guarantee superior performance; effective integration and task-appropriate division of labor are key to realizing the benefits of cooperation (Vaccaro, Almaatouq, and Malone, 2024).
Figure 1: Existing generative AI lacks proficiency in key aspects of co-construction of solutions to complex problems. We give a code synthesis example. (a) GPT-4 Turbo fails to follow preferences explicitly stated by the human expert. (b) Due to the lack of a persistent object representation, a modification request targeted toward one feature of the desired solution leads to the unwanted (and erroneous) modification of another feature. (c) The human expert modifies the generated code directly to remove inline assumptions and introduces general variables; such active participation is not demarcated and recorded by the AI and there is no facility to extract the implicit preferences and follow them elsewhere.
Scenarios that HAI-Co2 does not address. To more clearly delineate which class of scenarios we want HAI-Co2 to address, we now give some examples where HAI-Co2 is not a good fit for solving problems. (i) If the problem is simple (as opposed to a complex expert-domain problem), then AI models probably can solve it autonomously. (ii) Even many complex problems may be solvable autonomously by AI (now or in the future) if a full specification is available that can be checked automatically. Games like Go and chess are such examples in which objective co-construction is not necessary, i.e., a full specification of the objective can be easily obtained. (iii) Another class of scenarios may be best addressed by an expert managing the solution process and utilizing the agent for solving specific subtasks. Somebody working through their email inbox after coming back from vacation may want to closely oversee this process (e.g., not letting the agent send email autonomously), but may be happy to give specific subtasks to the agent (e.g., ask correspondents that requested reviews to re-request them).
In contrast to scenarios in which the cooperation is dominated by either the agent (ii) or the expert (iii), HAI-Co2 is intended for equal-partner scenarios in which the two work together as partners leveraging their complementary strengths.
Is the current state of generative AI enough? As a running example, we will use a code generation example that illustrates some of the bottlenecks of GPT-4 Turbo, an update [^1] We use gpt-4-1106-preview. of the original GPT-4 model (OpenAI, 2023). In Figure 1 (a), GPT-4 Turbo ignores the human expert’s explicit instructions to generate a complete Python code with the required module specifications, echoing Xie et al. (2024)’s observation that current language-based AI agents lack task adherence. After repeated prompting with partial code snippets, the process produces a complete – albeit faulty – code. This limitation is even more serious when more varied and realistic expressions of human preferences are taken into account – for the human expert to contribute productively, one must allow preferences expressed via explicit instructions, binary choice, ranking, etc. Current generative AI solutions do not facilitate such multi-modal preference incorporation. Figure 1 (b) shows unreliable debugging attempts. Specifically, the LLM performs unrelated (and faulty) edits to address a bug and even introduces new errors. This demonstrates that existing LLMs struggle with handling complex, modular software code (Jiang et al., 2024). The common practice is that the human (as a knowledgeable expert who keeps track of overall context) identifies faulty output and repeatedly prompts the model to guide it to the correct generation – this is implicitly adopting a co-construction paradigm. However, Figure 1 (c) shows that current modes of human-AI interaction cannot unlock the full potential of co-construction – direct modification of the co-constructed candidate solution by the human expert does not bear any special significance to the LLM, and it treats it as just another context. There is no explicit mechanism for the AI to learn implicit preferences expressed by the human through active participation.
While these examples are focused on code synthesis, there is evidence of similar shortcomings in other domains such as planning (Kambhampati et al., 2024), radiology (Lecler, Duron, and Soyer, 2023) and clinical decision making (Hager et al., 2024). Carroll et al. (2019), for example, demonstrate the necessity of incorporating explicit “human awareness” in a version of the collaborative game Overcooked, providing evidence that agents fail to coordinate with human subjects without such awareness. Code synthesis in particular – and the experience from day-to-day use of generative AI for solving complex problems in general – points toward co-construction as a naturally evolving problem solving paradigm where the human expert tries to search for the optimal solution by interacting with the AI. However, the current state of generative AI hinders its role as a reliable partner in successful co-construction. This is because the “one-directional” interaction between human and AI typical of how AI agents are used today often fails to steer the co-construction towards a solution that satisfies the user’s constraints.
In summary, prior work has laid out the inherent shortcomings of present-day generative AI for complex problem solving (see also Section 5 for a much more detailed discussion of prior work). This motivates our alternative emphasis on human-AI co-construction as a paradigm for solving complex problems in expert domains.
Ethical considerations. One aspect of HAI-Co2 carries considerable risk: the co-construction of the objective. This opens the door to manipulation by the agent. For example, LLMs that fail to solve a goal have been observed to redefine it to be easier Anthropic (2025a).
Influenceable reward functions have been studied by Carroll et al. (2024). They write: “We show that despite its convenience, the static-preference assumption may undermine the soundness of existing alignment techniques, leading them to implicitly reward AI systems for influencing user preferences in ways users may not truly want.” and: “…  suggesting that a straightforward solution to the problems of changing preferences may not exist.”
Even if there is no general solution to the problem of influenceable reward functions, it is possible to build in safeguards in the context of HAI-Co2. Specifically, if there is a conflict between user preferences and AI preferences, then we can mandate that user preferences prevail. This can be implemented by a “monitor” agent – a secondary agent that cannot be manipulated by the primary agent and is responsible for alerting the expert to objective changes that were not clearly communicated. Alternatively, we can employ the methodology of alignment to discourage unwanted manipulation of the objective by the agent. For example, we can devise a set of rules that should govern objective co-construction (e.g., “a change to the objective must be clearly communicated to the human”), create a synthetic dataset that embodies the rules and then train the agent on this dataset using supervised finetuning or reinforcement learning.
More generally, we believe that a paradigm of close cooperation is a promising approach to addressing many of the hard ethical problems that AI faces. If the agent takes an initial problem statement, goes off and comes back with a complete solution, then that means that the human cannot make any course corrections. This is true both for initial decisions (if she were part of the co-construction process, the human may realize that some initial decisions were based on wrong assumptions and correct them) and for decisions made by the agent (in which the human is not involved in autonomous problem solving and therefore cannot influence). Our co-construction paradigm ensures consistent human participation in shaping the solution as it is being constructed. Similarly, if the human is involved in the step-by-step co-construction of the solution, then she will have a good understanding of its inner workings and the motivation for its parts; thus, cooperation can be effective in bringing about some measure of explainability and (by extension) transparency with respect to ethical concerns.
In summary, making the objective influenceable by the agent is a risk. But there are promising solutions for addressing it (e.g., adding a monitoring agent for supervision). In addition, the close cooperation of human and agent in HAI-Co2– which in contrast to an autonomous approach ensures that the human is involved in all aspects of the co-construction process – is a form of artificial intelligence that addresses some ethics problems of AI by design, e.g., it supports transparency with respect to ethical concerns.
Our contribution. In this article, we take steps towards formalizing co-constructive problem solving and thereby aim to address important limitations of current generative AI models. In contrast to approaches in which problems are solved autonomously by AI or – conversely – in which AI is an assistant without autonomy that is limited to executing tasks clearly defined by the human, we view co-constructive problem solving as a process that involves the two parties as equal partners, each contributing complementary strengths. Concretely, we present HAI-Co2, a conceptual framework that facilitates human-AI co-construction. HAI-Co2 introduces multiple levels of abstraction to the candidate solution, providing a seamless interface for the human expert and the AI agent to modify and keep track of the complex, modular, co-constructed candidate solution. HAI-Co2 supports objective co-construction – i.e., both the solution and the objective are co-constructed by human and agent – by allowing multi-modal preference input from the human expert, with natural language as the central mediator to capture information-rich guidance signals, along with other forms of active expert participation, such as categorical choice-based preference. Solution co-construction in HAI-Co2 is conceived as search where the candidate solution (represented on multiple levels of abstraction) is iteratively revised to maximize the its utility, modeled by the preference model. While several components of HAI-Co2 have been explored in prior research independently across different domains (see Section 5), we are the first to bring them together under a unified conceptual umbrella and to show that, enabled by current generative AI, they have the potential to address the major challenges in solving complex expert-domain problems.
Figure 2: Illustration of the hierarchy of construction spaces in HAI-Co2. Each point ${\bm{x}}$ symbolizes a candidate solution (on a certain level of abstraction), e.g., a software program. The topology of the space is specified by a suitable neighborhood structure (as illustrated for point ${\bm{x}}$). Each point is associated with a latent utility $u^{t}$, possibly multi-dimensional and comprised of local utilities $u_{1}^{t},\ldots,u_{K}^{t}$, and preferential information (e.g., ${\bm{x}}\succ{\bm{x}}^{\prime}$: solution ${\bm{x}}$ is better than ${\bm{x}}^{\prime}$) that provides information about promising regions in the space. The relationship between the different abstraction levels is specified by the abstraction mappings $f_{j}$ resp. the (inverse) refinement mappings $f^{(-1)}_{j}$.
3 Towards a formalization of HAI-Co2